# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=false, bottomline=false, skipabove=skipbelow=false

### v1

@@ 6. Related Works
The SE community has extensively focused on mitigating unfairness in DNNs. However, SE literature on mitigating unfairness in LLMs is sparse. The approach followed by the Neufair (Dasu et al., 2024) and DICE (Monjezi et al., 2023) algorithms is related to attention head pruning for fairness in LLMs. DICE uses a fault localization technique via do logic to identify the top $k$ neurons negatively affecting fairness. Such an approach is used by the state-of-the-art FASP (Zayed et al., 2024) algorithm in which the effect of attention heads, instead of neurons, is analyzed. Our search space construction and solution using SA resembles the NeuFair algorithm. NeuFair (Dasu et al., 2024) mitigates unfairness in pre-trained deep neural networks (DNNs) by identifying a good set of neurons to drop. At each step of the SA search, it creates a copy of the DNN, drops a set of neurons, and analyzes their effect on fairness and model utility. Since it directly performs inference on the unfair DNNs, their algorithm does not scale to LLMs. The experimental analysis in the paper limits the SA search to small DNNs with at most 384 neurons and low inference times. Even if we apply such an approach for structured pruning at the attention head level instead of unstructured pruning of neurons in LLMs, it would be infeasible as just one round of inference during the state space exploration would take several minutes. While we quantitatively evaluate the state-of-the-art attention-pruning strategy for bias reduction, we briefly discuss other bias mitigation strategies in LLMs in three stages of pre-processing, in-processing, and post-processing:
@@
In-processing methods modify the model’s parameters and optimization process. Methods such as ADELE (Lauscher et al., 2021) introduce adapter layers for parameter-efficient fine-tuning to reduce bias. Regularization techniques can be applied to penalize the model for producing biased outputs (Mozafari et al., 2020). Additionally, debiasing word embeddings can help reduce bias in the model’s representation of words by reducing gendered or racial stereotypes (Gonen and Goldberg, 2019; Bolukbasi et al., 2016; Wang et al., 2020). Several loss functions, such as declustering loss (Garimella et al., 2021) and contrastive loss (Liu et al., 2021; He et al., 2022) are also shown to be effective for mitigating bias. Zhang et al. (2023) leveraged gradient-based explanations to target sensitive attributes and adjust training for balanced fairness and performance. Apart from that, Dhingra et al. (2023) used style transfer to replace harmful language while maintaining the original meaning.
Post-processing techniques modify the model’s inference behavior after complete training. This may involve altering the model before inference or the model output during/after inference. For instance, Gehman et al. (2020) proposed token-blocking methods during decoding to prevent the generation of harmful or biased terms. Hauzenberger et al. (2022) introduced sparse debiasing subnetworks that are trained separately and can be applied to the model at inference time. Qian et al. (2021) performed keyword-based distillation to remove bias during inference. Tokpo and Calders (2022) identify biased tokens and replace them with less stereotypical terms. Post-processing approaches are beneficial since dataset collection and LLM training are expensive processes that we may not always be able to repeat if issues of unfairness are found. Furthermore, post-processing mitigation approaches are sometimes the only viable option with the status quo of utilizing large pre-trained LLMs that take a huge amount of resources to train.
@@ 7. Discussion
Limitations Our metric relies on surrogate DNNs to approximate the bias and perplexity. Additionally, we use SA to explore the combinatorial search space. Both contribute to randomness in the search and our results can be sub-optimal. For example, in the GPT-Neo-125M LLM, we can only achieve better perplexity than FASP. We hypothesize this is due to the error in DNN predictions since they are not $100\%$ accurate. Additionally, while SA has statistical guarantees on convergence, it often requires an infeasible number of iterations to reach the optimal state. Another limitation of our approach is that the offline dataset creation to train surrogate DNNs may take time, depending on the LLM and dataset size. For example, the one-time offline data collection of $27,000$ samples and training of surrogate DNNs takes $\approx 1900$ GPU hours in total across multiple GPUs.
@@
Threat to validity To address the internal validity and ensure our finding does not lead to an invalid conclusion, we follow the established SE guidelines by repeating the experiments multiple times and reporting the mean and standard deviation. One challenge is the architecture of DNNs, where we manually performed hyperparameter tuning to find DNNs with ideal accuracy. The results might vary with different architectures as the SA search relies on them. We also observed outliers in some experiments. For example, in Table 5, we see that the bias of Llama-2-7B is higher when only the bias DNN is used in the cost function compared to the linear combination of both DNNs. This phenomenon was observed after repeating the experiment multiple times. We hypothesize that the relation between bias and perplexity is complex for Llama-2-7B, leaving more analysis to future work. To ensure our results are generalizable, we perform experiments on six different LLMs of varying sizes. Due to limited computational resources, we limit our analysis to the HolisticBias metric to measure bias and the WikiText-2 dataset to measure perplexity. Although the state-of-the-art method used a similar experimental setup, it is unknown whether our method will generalize to all LLMs and bias metrics.
@@ 8. Conclusion
In this paper, we tackle the problem of mitigating unfairness in pre-trained LLMs by carefully pruning attention heads. While the attention head pruning is a computationally hard problem, we found that techniques based on surrogate simulated annealing are effective and scalable to improve the fairness of LLMs while preserving the utility. We implemented our approach, Attention Pruning, and showed that it can identify an ideal subset of attention heads that disparately contribute to unfairness and outperform the state-of-the-art post-processing bias mitigator.
For future work, one promising direction is to build a genetic algorithm to explore the top $k$ states found during the search instead of choosing the best state. Another direction is to apply the pruning idea to improve other Responsible AI requirements, such as privacy and robustness.

### v2

@@ 6. Related Works
The SE community has extensively focused on mitigating unfairness in DNNs, and the literature on mitigating unfairness in LLMs is sparse. DICE (Monjezi et al., 2023) uses a fault localization technique via do logic to identify the top $k$ neurons negatively affecting fairness. FASP uses a similar approach to analyze the effect of pruning attention heads instead of neurons. Our search space construction and solution using SA resembles the NeuFair algorithm. NeuFair (Dasu et al., 2024) mitigates unfairness in pre-trained DNNs by identifying a good set of neurons to drop. At each step of the SA search, it creates a copy of the DNN, drops a set of neurons, and analyzes their effect on fairness and model utility. Since it directly performs inference on the unfair DNN, it does not scale to LLMs. The experimental analysis in the paper limits the SA search to small DNNs with at most 384 neurons and low inference times. Even if we apply such an approach for structured pruning at the attention head level instead of unstructured pruning of neurons in LLMs, it would be infeasible, as just one round of inference during the state space exploration would take several minutes. While we quantitatively evaluate against the state-of-the-art attention-pruning strategy for bias reduction, we briefly discuss other bias mitigation strategies in LLMs in three stages of pre-processing, in-processing, and post-processing:
@@
In-processing methods modify the model’s parameters and optimization process. ADELE (Lauscher et al., 2021) uses adapter layers for parameter-efficient fine-tuning to reduce bias. Regularization techniques can be applied to penalize the model for producing biased outputs (Mozafari et al., 2020). Debiasing word embeddings can help reduce bias in the model’s representation of words by reducing gendered or racial stereotypes (Gonen and Goldberg, 2019; Bolukbasi et al., 2016; Wang et al., 2020). Alternate loss functions such as declustering loss (Garimella et al., 2021) and contrastive loss (Liu et al., 2021; He et al., 2022) have also been shown to be effective for mitigating bias. Zhang et al. (2023) leveraged gradient-based explanations to target sensitive attributes and adjust training for balanced fairness and performance.Dhingra et al. (2023) used style transfer to replace harmful language while maintaining the original meaning.
Post-processing techniques modify the model’s inference behavior after complete training. Gehman et al. (2020) proposed token-blocking methods during decoding to prevent the generation of harmful or biased terms. Hauzenberger et al. (2022) introduced sparse debiasing subnetworks that are trained separately and can be applied to the model at inference time. Qian et al. (2021) performed keyword-based distillation to remove bias during inference. Tokpo and Calders (2022) identify biased tokens and replace them with less stereotypical terms. Post-processing approaches are beneficial since dataset collection and training are expensive processes that we may not be able to repeat if fairness issues are found. Furthermore, post-processing mitigation approaches are sometimes the only viable option with the status quo of utilizing large pre-trained LLMs that take a huge amount of resources to train.
@@ 7. Discussion
Qualitative Examples To better understand how fairness-aware pruning strategies minimize bias, we provide a qualitative example in Table 9. We analyze two prompts belonging to the sex and cisgender subgroups of the Gender group in HolisticBias. We sample 40 tokens from the LLM or until the STOP token is generated, whichever comes first. The responses in all three cases may seem incoherent or incomplete since only 40 tokens are sampled. From the responses, we see that the baseline LLM is biased against the sex subgroup as it generates more toxic content from the prompt compared to the cisgender prompt. FASP and AP do not guarantee zero toxicity. They reduce bias by minimizing the toxicity difference, thereby ensuring similar toxicity across all subgroups. In general, from Table 6, AP is better than other methods at reducing bias across all $33,789$ test set prompts and 8 subgroups in the Gender group of HolisticBias.
Practical Significance of Bias Reduction From Table 6, we see that AP usually has better bias than FASP by $\approx 0.02$ and better perplexity up to $\approx 6$ points or up to $\approx 11\%$. To ensure that this is practically significant, we perform the Cliff’s $\delta$ and Mann-Whitney U statistical tests. We repeat the experiment for 25 random seeds each for AP and FASP. For GPT-J-6B, we observe a $\delta=-0.968$ (95% CI $\approx[-0.994,-0.847]$), indicating a large effect; $\delta$ is negative as lower bias is better. For the Mann-Whitney test, we get $U=10.0$ and (one-sided) $p=2.319\times 10^{-9}$, indicating a 1 in $\approx 400$ million chance that FASP is similar to AP, thus rejecting the null hypothesis that both are from the same distribution. Similarly, for Llama-2-7B, we observe a $\delta=-0.9392$ (95% CI $\approx[-0.984,-0.785]$) which indicates a large effect. For the Mann-Whitney test, we get $U=19$ and (one-sided) $p=6.539\times 10^{-9}$. The perplexity is seed-independent and only depends on the WikiText-2 test split. Apart from achieving lower bias, AP also achieves lower perplexity. State-of-the-art techniques in literature claim a reduction of $\approx 1-2$ points (Sun et al., 2025; Dai et al., 2019) and up to 10% reduction in perplexity (Maini et al., 2024) as significant.
Transferability of Surrogate DNNs While the idea of utilizing surrogate DNNs to speed up the SA search is general, the exact architecture of the DNNs defined in Table 2 is not intended to be transferable to different datasets and fairness metrics. The size and number of layers used in the DNNs depend on the datasets $\mathcal{D}_{\text{bias}}$ and $\mathcal{D}_{\text{ppl}}$ in Algorithm 1 used to train the DNNs. We determine the architecture and hyperparameters using a hyperparameter search to minimize the validation dataset loss. The standard data preprocessing steps outlined in Section 5 can be applied to all datasets to minimize variance. Furthermore, our experiments with the Race attribute (Table 8) demonstrate that the concept of surrogate DNNs extends beyond gender bias.
@@
Limitations We rely on surrogate DNNs to estimate the bias and perplexity. Additionally, we use SA to explore the combinatorial search space. Both contribute to randomness in the search, and our results may be sub-optimal. For example, for GPT-Neo-125M, we can only achieve better perplexity than FASP. We hypothesize this is due to the error in DNN predictions since they are not $100\%$ accurate. Additionally, while SA has statistical guarantees on convergence, it often requires an infeasible number of iterations to reach the optimal state. Another limitation of our approach is that the offline dataset creation to train surrogate DNNs may take time, depending on the LLM and dataset size. For example, the one-time offline data collection of $27,000$ samples and training of surrogate DNNs takes $\approx 1900$ GPU hours in total across multiple GPUs for Llama-2-7B.
@@
Threat to validity To address the internal validity and ensure our finding does not lead to an invalid conclusion, we follow the established SE guidelines by repeating the experiments multiple times and reporting the mean and standard deviation. One challenge is the architecture of DNNs, where we manually performed hyperparameter tuning to find DNNs with ideal accuracy. The results might vary with different architectures, as the SA search relies on them. We also observed outliers in some experiments. For example, in Table 5, we see that the bias of Llama-2-7B is higher when only the bias DNN is used in the cost function compared to the linear combination of both DNNs. This phenomenon was observed after repeating the experiment multiple times. We hypothesize that the relation between bias and perplexity is complex for Llama-2-7B, leaving more analysis to future work. To ensure that our results are generalizable, we performed experiments on six different LLMs of varying sizes. Due to limited computational resources, we only study the HolisticBias and WikiText-2 datasets. Although the state-of-the-art method used a similar experimental setup, it is unknown whether our method will generalize to all LLMs and bias metrics.
@@ 8. Conclusion
In this paper, we tackle the problem of mitigating unfairness in pre-trained LLMs by carefully pruning attention heads. Although attention head pruning is a computationally hard problem, we found that techniques based on surrogate simulated annealing are effective and scalable to improve the fairness of LLMs while preserving the utility. We implemented our approach, Attention Pruning, and showed that it outperforms the state-of-the-art post-processing bias mitigator by identifying an ideal subset of heads to prune.
For future work, developing genetic algorithms that explores the top $k$ states found during the search, rather than selecting the best state, is a promising direction. Another direction is to extend the head pruning idea to other responsible AI requirements, such as privacy and robustness.